# RIME: Robust Preference-based Reinforcement Learning with Noisy Preferences

### v2

@@ Appendix A RIME Algorithm Details
3:// Unsupervised Pre-training
4:for each pre-training step $t$ do
5:     Collect $\mathbf{s}_{t+1}$ by taking $\mathbf{a}_{t}\sim\pi_{\phi}(\mathbf{a}_{t}|\mathbf{s}_{t})$
6:     Compute normalized intrinsic reward $r^{\text{int}}_{\text{norm},t}\leftarrow r^{\text{int}}_{\text{norm}}(\mathbf{s}_{t})$ as in Equation (9)
7:     Store transitions $\mathcal{B}\leftarrow\mathcal{B}\cup\left\{(\mathbf{s}_{t},\mathbf{a}_{t},\mathbf{s}_{t+1},r^{\text{int}}_{\text{norm},t})\right\}$
8:     for each gradient step do
9:         Sample minibatch $\left\{(\mathbf{s}_{j},\mathbf{a}_{j},\mathbf{s}_{j+1},r^{\text{int}}_{\text{norm},j})\right\}^{B}_{j=1}\sim\mathcal{B}$
10:         Optimize policy and Q-network with respect to $\phi$ and $\theta$ using SAC
11:         // Warm start
12:         Update reward model $\hat{r}_{\psi}$ according to Equation (10)
13:     end for
14:end for
15:// Online training
16:Initialize the maximum KL divergence value $\rho=\infty$
17:Initialize a dataset of noisy preferences $\mathcal{D}_{\text{noisy}}\leftarrow\emptyset$
18:for each training step $t$ do
19:     // Robust reward learning
20:     if $t\%K==0$ then
21:         Generate queries from replay buffer $\{(\sigma^{0}_{i},\sigma^{1}_{i})\}_{i=1}^{N_{\text{query}}}\sim\mathcal{B}$ and corresponding human feedback $\{\tilde{y}_{i}\}_{i=1}^{N_{\text{query}}}$
22:         Store preferences $\mathcal{D}_{\text{noisy}}\leftarrow\mathcal{D}_{\text{noisy}}\cup\{(\sigma^{0}_{i},\sigma^{1}_{i},\tilde{y}_{i})\}_{i=1}^{N_{\text{query}}}$
23:         Compute lower bound $\tau_{\text{lower}}$ according to Equation (5)
24:         Filter trustworthy samples $\mathcal{D}_{t}$ using lower bound $\tau_{\text{lower}}$ as in Equation (6)
25:         Flip labels using upper bound $\tau_{\text{upper}}$ to obtain dataset $\mathcal{D}_{f}$ as in Equation (7)
26:         Update reward model $\hat{r}_{\psi}$ with samples from $\mathcal{D}_{t}\cup\mathcal{D}_{f}$ according to Equation (8)
27:         Relabel entire replay buffer $\mathcal{B}$ using $\hat{r}_{\psi}$
28:         Update parameter $\rho$ with the maximum KL divergence between predicted and annotated labels in dataset $\mathcal{D}_{t}\cup\mathcal{D}_{f}$
29:     end if
30:     for each timestep $t$ do
31:         Collect $\mathbf{s}_{t+1}$ by taking $\mathbf{a}_{t}\sim\pi_{\phi}(\mathbf{a}_{t}|\mathbf{s}_{t})$
32:         Store transitions $\mathcal{B}\leftarrow\mathcal{B}\cup\{(\mathbf{s}_{t},\mathbf{a}_{t},\mathbf{s}_{t+1},\widehat{r}_{\psi}(\mathbf{s}_{t},\mathbf{a}_{t}))\}$
33:     end for
34:     for each gradient step do
35:         Sample minibatch from replay buffer $\{(\mathbf{s}_{j},\mathbf{a}_{j},\mathbf{s}_{j+1},\hat{r}_{\psi}(\mathbf{s}_{j},\mathbf{a}_{j})\}_{j=1}^{B}\sim\mathcal{B}$
36:         Optimize policy and Q-network with respect to $\phi$ and $\theta$ using SAC
37:     end for
38:end for
@@ Appendix B Effects of biased reward model
Previous work empirically showed the detrimental impact of noisy preference on the reward model (Lee et al., 2021a). To further demonstrated the effects of a biased reward model, we introduce the following theorem and give the proof as follows.
@@
Appendix C Proofs for Theorem 4.1
Theorem C.1.
@@
$D_{\mathrm{KL}}\left(P_{\psi}(x)\|\tilde{y}(x)\right)\geq-\ln\rho+\frac{\rho}{2}+\mathcal{O}(\rho^{2})$ (22)
@@
    Then if the label is corrupted, denoted by $\tilde{y}_{c}$ (i.e.$\tilde{y}_{c}=(1,0)$ in this case), the KL divergence between predicted label and corrupted label is formulated as follows:
    $D_{\mathrm{KL}}(P_{\psi}\|\tilde{y}_{c})=-\ln P_{\psi}(0)\geq-\ln(1-\exp(-\rho))$ (25)
@@
    Then if the label is corrupted, i.e.$\tilde{y}_{c}\in\{(0,1),(1,0)\}$, the KL divergence between predicted label and corrupted label is formulated as follows:
    $D_{\mathrm{KL}}(P_{\psi}\|\tilde{y}_{c})\geq$ $\min(-\ln P_{\psi}(0),-\ln(1-P_{\psi}(0)))=-\ln p$ (30)
@@
$D_{\mathrm{KL}}(P_{\psi}\|\tilde{y}_{c})\geq-\ln(1-\exp(-\rho))=-\ln\rho+\frac{\rho}{2}+\mathcal{O}(\rho^{2})$ (37)
@@ D.2 Implementation Details
For the implementation of baselines, we use their corresponding publicly released repositories that are shown in Table 6. SAC serves as a performance upper bound, because it uses a ground-truth reward function which is unavailable in PbRL settings for training. The detailed hyperparameters of SAC are shown in Table 7. PEBBLE’s settings remain consistent with its original implementation, and the specifics are detailed in Table 8. For SURF, RUNE, MRN, and RIME, most hyperparameters are the same as those of PEBBLE and other hyperparameters are detailed in Table 9, 10, 11, and 12, respectively. The total amount of feedback and feedback amount per session in each experimental condition are detailed in Table 1. The reward model comprises an ensemble of three MLPs. Each MLP consists of three layers with 256 hidden units, and the output of the reward model is constrained using the tanh activation function.
@@ D.3 Details of scripted teachers
For a pair of trajectory segments $(\sigma^{0},\sigma^{1})$ with length $H$, where $\sigma^{i}=\{(\mathbf{s}_{1}^{i},\mathbf{a}^{i}_{1}),\dots,(\mathbf{s}_{H}^{i},\mathbf{a}_{H}^{i})\},(i=0,1)$. The ground-truth reward function from environment is $r(\mathbf{s},\mathbf{a})$. Then scripted teachers are defined as follows (Lee et al., 2021a):
@@
Equal: Equal teacher gives equal preference $(0.5,0.5)$ if the difference between the cumulative rewards of two segments are small. In particular, if $|\sum_{t=1}^{H}r(\mathbf{s}_{t}^{1},\mathbf{a}_{t}^{1})-\sum_{t=1}^{H}r(\mathbf{s}_{t}^{0},\mathbf{a}_{t}^{0})|<\delta$, then it return the label $(0.5,0.5)$. $\delta=\frac{H}{T}R_{\text{avg}}\epsilon_{\text{adapt}}$, where T is the episode length, $R_{\text{avg}}$ is the average returns of current policy, $\epsilon_{\text{adapt}}$ is a hyperparameters and is set to 0.1 in experiments, following the setting in Lee et al. (2021a).
@@ D.4 Details of experiments with human teachers
Human experiments adopt an online paradigm consistent with PEBBLE’s pipeline, where agent training alternates with reward model training. When it is the timestep to collect preferences (post-agent training and pre-reward training), the training program generates segment pairs , saving each segment in GIF format. The program then pauses, awaiting the input of human preferences. At this juncture, human labelers engage, reading the paired segments in GIF format and labelling their preferences. Subsequently, this annotated data is fed into the training program for reward model training.
To ensure a fair comparison between RIME and PEBBLE, for one human labeler, we start the training programs for RIME and for PEBBLE simultaneously. When both training programs are waiting for inputting preferences, we collate all GIF pairs from both RIME and PEBBLE and shuffle the order. Then the human labeler starts working. Therefore, from the labeler’s perspective, he/she does not know which algorithm the currently labeled segment pair comes from and just focuses on labelling according to his/her preference. The labeled data is then automatically directed to the respective training programs. We conduct this experiment parallelly on each of the five labelers, thus preferences from different users do not get mixed.
@@ Appendix E Additional Experiment Results
Effects of hyperparameters of RIME. We investigate how the hyperparameters of RIME affect the performance under noisy feedback settings. In Figure 7 we plot the learning curves of RIME with different set of hyperparameters: (a) coefficient $\alpha$ in the lower bound $\tau_{\text{lower}}$: $\alpha\in\{0.3,0.4,0.5,0.6\}$, (b) maximum value of $\beta_{t}$: $\beta_{\max}\in\{1,3,5,10\}$, (c) decay rate $k\in\{0.01,1/30,0.06,0.1\}$, and (d) upper bound of KL divergence $\tau_{\text{upper}}\in\{2\ln(10),3\ln(10),4\ln(10)\}$.
For the coefficient $\alpha$ in the lower bound $\tau_{\text{lower}}$, we find the theoretical value $\alpha=0.5$ performs the best. The maximum weight $\beta_{\max}$ and decay rate $k$ control the weight of uncertainty term in the lower bound $\tau_{\text{lower}}$: $\beta_{t}=\max(\beta_{\min},\beta_{\max}-kt)$. The combination of $\beta_{\max}=3$ and $k=1/30$ also performs optimally. Due to the quite limited feedback amount (1000 feedback) and training epochs for the reward model (around $150\sim 200$ epochs on Walker-walk), RIME is sensitive to the weight of uncertainty term. If ones try to increase $\beta_{\max}$ to add more tolerance for trustworthy samples in early-stage, we recommend to increase the decay rate $k$ simultaneously so that the value of $\beta_{t}$ decays to its minimum within about $1/3$ to $1/2$ of the total epochs. For the upper bound $\tau_{\text{upper}}$, although we use $3\ln(10)$ for balanced performance on DMControl tasks, individually fine-tuning $\tau_{\text{upper}}$ can further improve the performance of RIME on the corresponding task, such as using $\tau_{\text{upper}}=4\ln(10)$ for Walker-walk.
@@
Figure 8: Uncertainty term analysis on Walker-walk using 1000 feedback with $\epsilon=0.3$, across ten runs.
@@
For reference, we also include a group devoid of any uncertainty term, termed the “None” group. As shown in Figure 8, the KL metric outperforms the other approaches on Walker-walk with an error rate of $\epsilon=0.3$. This might be because the disagreement metric fluctuates violently at every query times, often leading to excessive trust in new data, which hinders the stabilization of the lower bound.

### v3

@@ Appendix A RIME Algorithm Details
3:for each pre-training step $t$ do $\triangleright$ Unsupervised Pre-training
4:     Collect $\mathbf{s}_{t+1}$ by taking $\mathbf{a}_{t}\sim\pi_{\phi}(\mathbf{a}_{t}|\mathbf{s}_{t})$
5:     Compute normalized intrinsic reward $r^{\text{int}}_{\text{norm},t}\leftarrow r^{\text{int}}_{\text{norm}}(\mathbf{s}_{t})$ as in Equation (9)
6:     Store transitions $\mathcal{B}\leftarrow\mathcal{B}\cup\left\{(\mathbf{s}_{t},\mathbf{a}_{t},\mathbf{s}_{t+1},r^{\text{int}}_{\text{norm},t})\right\}$
7:     for each gradient step do
8:         Sample minibatch $\left\{(\mathbf{s}_{j},\mathbf{a}_{j},\mathbf{s}_{j+1},r^{\text{int}}_{\text{norm},j})\right\}^{B}_{j=1}\sim\mathcal{B}$
9:         Optimize policy and Q-network with respect to $\phi$ and $\theta$ using SAC
10:         Update reward model $\hat{r}_{\psi}$ according to Equation (10) $\triangleright$ Warm start
11:     end for
12:end for
13:Initialize the maximum KL divergence value $\rho=\infty$
14:Initialize a dataset of noisy preferences $\mathcal{D}_{\text{noisy}}\leftarrow\emptyset$
15:for each training step $t$ do $\triangleright$ Online training
16:     if step to query preferences then $\triangleright$ Robust reward learning
17:         Generate queries from replay buffer $\{(\sigma^{0}_{i},\sigma^{1}_{i})\}_{i=1}^{N_{\text{query}}}\sim\mathcal{B}$ and corresponding human feedback $\{\tilde{y}_{i}\}_{i=1}^{N_{\text{query}}}$
18:         Store preferences $\mathcal{D}_{\text{noisy}}\leftarrow\mathcal{D}_{\text{noisy}}\cup\{(\sigma^{0}_{i},\sigma^{1}_{i},\tilde{y}_{i})\}_{i=1}^{N_{\text{query}}}$
19:         Compute lower bound $\tau_{\text{lower}}$ according to Equation (5)
20:         Filter trustworthy samples $\mathcal{D}_{t}$ using lower bound $\tau_{\text{lower}}$ as in Equation (6)
21:         Flip labels using upper bound $\tau_{\text{upper}}$ to obtain dataset $\mathcal{D}_{f}$ as in Equation (7)
22:         Update reward model $\hat{r}_{\psi}$ with samples from $\mathcal{D}_{t}\cup\mathcal{D}_{f}$ according to Equation (8)
23:         Relabel entire replay buffer $\mathcal{B}$ using $\hat{r}_{\psi}$
24:         Update parameter $\rho$ with the maximum KL divergence between predicted and annotated labels in dataset $\mathcal{D}_{t}\cup\mathcal{D}_{f}$
25:     end if
26:     Collect $\mathbf{s}_{t+1}$ by taking $\mathbf{a}_{t}\sim\pi_{\phi}(\mathbf{a}_{t}|\mathbf{s}_{t})$
27:     Store transitions $\mathcal{B}\leftarrow\mathcal{B}\cup\{(\mathbf{s}_{t},\mathbf{a}_{t},\mathbf{s}_{t+1},\widehat{r}_{\psi}(\mathbf{s}_{t},\mathbf{a}_{t}))\}$
28:     for each gradient step do
29:         Sample minibatch from replay buffer $\{(\mathbf{s}_{j},\mathbf{a}_{j},\mathbf{s}_{j+1},\hat{r}_{\psi}(\mathbf{s}_{j},\mathbf{a}_{j})\}_{j=1}^{B}\sim\mathcal{B}$
30:         Optimize policy and Q-network with respect to $\phi$ and $\theta$ using SAC
31:     end for
32:end for
@@ Appendix B Effects of biased reward model
Previous work empirically showed the detrimental impact of noisy preference on the reward model (Lee et al., 2021a). To further demonstrate the effects of a biased reward model, we introduce the following theorem and give the proof as follows.
@@
Appendix C Proof for Theorem 4
Theorem (4).
@@ Appendix C Proof for Theorem 4
$D_{\mathrm{KL}}\left(\tilde{y}(x)\|P_{\psi}(x)\right)\geq-\ln\rho+\frac{\rho}{2}+\mathcal{O}(\rho^{2})$ (22)
@@
    Then if the label is corrupted, denoted by $\tilde{y}_{c}$ (i.e.,$\tilde{y}_{c}=(1,0)$ in this case), the KL divergence between the predicted label and the corrupted label is formulated as follows:
    $D_{\mathrm{KL}}(\tilde{y}_{c}\|P_{\psi})=-\ln P_{\psi}(0)\geq-\ln(1-\exp(-\rho))$ (25)
@@
    Then if the label is corrupted, i.e.,$\tilde{y}_{c}\in\{(0,1),(1,0)\}$, the KL divergence between the predicted label and the corrupted label is formulated as follows:
    $D_{\mathrm{KL}}(\tilde{y}_{c}\|P_{\psi})\geq$ $\min(-\ln P_{\psi}(0),-\ln(1-P_{\psi}(0)))=-\ln p$ (30)
@@
$D_{\mathrm{KL}}(\tilde{y}_{c}\|P_{\psi})\geq-\ln(1-\exp(-\rho))=-\ln\rho+\frac{\rho}{2}+\mathcal{O}(\rho^{2})$ (37)
@@ D.2 Implementation Details
For the implementation of baselines, we use their corresponding publicly released repositories that are shown in Table 8. SAC serves as a performance upper bound because it uses a ground-truth reward function which is unavailable in PbRL settings for training. The detailed hyperparameters of SAC are shown in Table 9. PEBBLE’s settings remain consistent with its original implementation, and the specifics are detailed in Table 10. For SURF, RUNE, MRN, and RIME, most hyperparameters are the same as those of PEBBLE and other hyperparameters are detailed in Table 11, 12, 13, and 14, respectively. The total amount of feedback and feedback amount per session in each experimental condition are detailed in Table 1. The reward model comprises an ensemble of three MLPs. Each MLP consists of three layers with 256 hidden units, and the output of the reward model is constrained using the tanh activation function.
@@ D.3 Details of scripted teachers
For a pair of trajectory segments $(\sigma^{0},\sigma^{1})$ with length $H$, where $\sigma^{i}=\{(\mathbf{s}_{1}^{i},\mathbf{a}^{i}_{1}),\dots,(\mathbf{s}_{H}^{i},\mathbf{a}_{H}^{i})\},(i=0,1)$. The ground-truth reward function from the environment is $r(\mathbf{s},\mathbf{a})$. Then scripted teachers are defined as follows (Lee et al., 2021a):
@@
Equal: Equal teacher gives equal preference $(0.5,0.5)$ if the difference between the cumulative rewards of two segments is small. In particular, if $|\sum_{t=1}^{H}r(\mathbf{s}_{t}^{1},\mathbf{a}_{t}^{1})-\sum_{t=1}^{H}r(\mathbf{s}_{t}^{0},\mathbf{a}_{t}^{0})|<\delta$, then it return the label $(0.5,0.5)$. $\delta=\frac{H}{T}R_{\text{avg}}\epsilon_{\text{adapt}}$, where T is the episode length, $R_{\text{avg}}$ is the average returns of current policy, $\epsilon_{\text{adapt}}$ is a hyperparameter and is set to 0.1 in experiments, following the setting in Lee et al. (2021a).
@@ D.4 Details of experiments with human teachers
Human experiments adopt an online paradigm consistent with PEBBLE’s pipeline, where agent training alternates with reward model training. When it is the timestep to collect preferences (post-agent training and pre-reward training), the training program generates segment pairs, saving each segment in GIF format. The program then pauses, awaiting the input of human preferences. At this juncture, human labelers engage, reading the paired segments in GIF format and labeling their preferences. Subsequently, this annotated data is fed into the training program for reward model training.
To ensure a fair comparison between RIME and PEBBLE, for one human labeler, we start the training programs for RIME and PEBBLE simultaneously. When both training programs are waiting for inputting preferences, we collate all GIF pairs from both RIME and PEBBLE and shuffle the order. Then the human labeler starts working. Therefore, from the labeler’s perspective, he/she does not know which algorithm the currently labeled segment pair comes from and just focuses on labeling according to his/her preference. The labeled data is then automatically directed to the respective training programs. We conduct this experiment parallelly on each of the five labelers, thus preferences from different users do not get mixed.
@@ Appendix E Additional Experiment Results
Effects of hyperparameters of RIME. We investigate how the hyperparameters of RIME affect the performance under noisy feedback settings. In Figure 7 we plot the learning curves of RIME with a different set of hyperparameters: (a) coefficient $\alpha$ in the lower bound $\tau_{\text{lower}}$: $\alpha\in\{0.3,0.4,0.5,0.6\}$, (b) maximum value of $\beta_{t}$: $\beta_{\max}\in\{1,3,5,10\}$, (c) decay rate $k\in\{0.01,1/30,0.06,0.1\}$, and (d) upper bound of KL divergence $\tau_{\text{upper}}\in\{2\ln(10),3\ln(10),4\ln(10)\}$.
For the coefficient $\alpha$ in the lower bound $\tau_{\text{lower}}$, we find the theoretical value $\alpha=0.5$ performs the best. The maximum weight $\beta_{\max}$ and decay rate $k$ control the weight of uncertainty term in the lower bound $\tau_{\text{lower}}$: $\beta_{t}=\max(\beta_{\min},\beta_{\max}-kt)$. The combination of $\beta_{\max}=3$ and $k=1/30$ also performs optimally. Due to the quite limited feedback amount (1000 feedback) and training epochs for the reward model (around $150\sim 200$ epochs on Walker-walk), RIME is sensitive to the weight of uncertainty term. If one tries to increase $\beta_{\max}$ to add more tolerance for trustworthy samples in early-stage, we recommend increasing the decay rate $k$ simultaneously so that the value of $\beta_{t}$ decays to its minimum within about $1/3$ to $1/2$ of the total epochs. For the upper bound $\tau_{\text{upper}}$, although we use $3\ln(10)$ for balanced performance on DMControl tasks, individually fine-tuning $\tau_{\text{upper}}$ can further improve the performance of RIME on the corresponding task, such as using $\tau_{\text{upper}}=4\ln(10)$ for Walker-walk.
@@
Figure 8: Uncertainty term analysis on Walker-walk using 1000 feedback with error rate $\epsilon=0.3$, across ten runs.
@@
For reference, we also include a group devoid of any uncertainty term, termed the “None” group. As shown in Figure 8, the KL metric outperforms the other approaches on Walker-walk with an error rate of $\epsilon=0.3$. This might be because the disagreement metric fluctuates violently at every query time, often leading to excessive trust in new data, which hinders the stabilization of the lower bound.
Comparison with fixed lower bound. We conducted experiments to compare fixed lower bound with dynamic lower bound (RIME) and presented the results in Table 15. Notice that the convergence value of $\tau_{\text{lower}}$ in RIME are 0.972 and 0.711 for Walker and Button-press, respectively. Table 15 indicates that the dynamic lower bound employed by RIME outperforms the fixed lower bound method substantially. This superiority stems from RIME’s ability to adapt its lower bound value according to the situation during training. By contrast, employing a fixed lower bound might exacerbate incorrect selections either in the early or late phase of training, depending on whether the lower bound is small or large respectively. The issue of incorrect selection will in turn lead to cumulative errors and compromise the effectiveness of selection-based robust training methods.
Table 15: Ablation study of the lower bound on Walker and Button-press over 5 runs.
Method Value of $\tau_{\text{lower}}$ Walker ($\epsilon=0.3$) Button-press ($\epsilon=0.3$)
0.5 98
$\pm$ 92
36.8
$\pm$ 45.4
Fixed 0.7 179
$\pm$ 165
64.8
$\pm$ 26.4
lower 0.9 256
$\pm$ 88
58.0
$\pm$ 41.9
bound 1.1 468
$\pm$ 186
49.0
$\pm$ 37.6
1.3 325
$\pm$ 79
56.6
$\pm$ 34.0
RIME dynamic 741 $\pm$ 139 80.0 $\pm$ 27.7
Explore the error rate limits of RIME. To understand the boundaries of RIME, we conduct supplementary experiments in both the Walker-walk and Cheetah-run environments, varying the error rates $\epsilon$ within the range $\{35\%,40\%,45\%,50\%\}$, across 5 runs. The results, presented in Table 16, reveal that RIME failed with 45% and 40% noisy data in the Walker-walk and Cheetah-run environments, respectively. Interestingly, even with the feedback amount increased to ten times the minimum, it has not change the result of failure. The performance gains from increasing the amount of feedback are limited under high error rates ($\epsilon\geq 0.4$).
Table 16: Performance of RIME with different noise levels over 5 runs.
Environment Feedback Error rate
volume 35% 40% 45% 50%
Walker 1000 646.58 497.63 164.31 /
5000 / 500.64 210.86 145.05
10000 / 554.69 312.61 217.34
Cheetah 1000 403.72 246.51 / /
5000 440.4 347.4 / /
10000 503.57 393.0 / /
[…]
Table 17: Results of PEBBLE as the error rate increases with constant amount of feedback, across 5 runs.
Environment Feedback Error rate
volume 0.1 0.2 0.3
Walker 500 749
$\pm$ 123
490
$\pm$ 252
230
$\pm$ 172
Quadruped 2000 292
$\pm$ 166
171
$\pm$ 26
125
$\pm$ 38
Button-press 10000 93.1
$\pm$ 10.6
21.6
$\pm$ 15.3
17.8
$\pm$ 25.2
Hammer 20000 36.6
$\pm$ 41.4
20.0
$\pm$ 17.8
15.7
$\pm$ 12.0
Table 18: Results of PEBBLE as the feedback volume increases with constant error rate, across 5 runs. $N$ refers to the minimal feedback volume for each environment shown in Table 1.
Domain Environment Error rate Feedback volume
[…]
DMControl Walker 0.2 490
$\pm$ 252
656
$\pm$ 158
0.3 230
$\pm$ 172
431
$\pm$ 157
Quadruped 0.2 171
$\pm$ 26
212
$\pm$ 47
0.3 125
$\pm$ 38
165
$\pm$ 35
Meta-world Button-press 0.2 21.6
$\pm$ 15.3
26.2
$\pm$ 35.7
0.3 17.8
$\pm$ 25.2
22.0
$\pm$ 13.8